OpenStudy (anonymous):

Which of the following is the equation of the axis of symmetry of the quadratic equation y = 2x^2 + 24x + 62? a) -10 b) -6 c) 6 d) 10

OpenStudy (ybarrap):

2(x^2 + 12x + 32) = 2*(x + 6)^2 - 6^2 + 32 minimum occurs at x = -6, which is also the axis of symmetry. So axis of symmetry is x = -6 Ans. b) -6.